# Strong Screening Rules for Group-based SLOPE Models

### v1

@@ B.1 Penalty weights
The penalty weights for SGS provide variable and group FDR-control simultaneously, under orthogonal designs [14]. They are given by (where the indexing corresponds to the sorted variables/groups)
$v_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\alpha}F_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right)-\frac{1}{3\alpha}(1-\alpha)a_{j}w_{j}\right\},\;i=1,\dots,p,$ (11)
$w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{F_{\text{FN}}^{-1}(1-\frac{q_{g}i}{m})-\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}}{(1-\alpha)p_{j}}\right\},\;i=1,\dots,m,$ (12)
where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom, $F_{\mathcal{N}}$ is the cumulative distribution function of a folded Gaussian distribution, and $a_{j}$ is a quantity that requires estimation. The estimator $\hat{a}_{j}=\lfloor\alpha p_{j}\rfloor$ is proposed in [14]. As with gSLOPE (Appendix A.1), a relaxation is possible, giving the weights
@@
The mean weights defined in Equations 13 and 14 are used for all SGS numerical simulations in this manuscript (shown in Figure A2).
@@ B.2 Derivation of soft thresholding operator
    $\nabla_{i}f(\beta)\in[-\lambda\alpha v_{i},\lambda\alpha v_{i}]$: choose $y_{i}=\nabla_{i}f(\beta)/\alpha\lambda$.
@@
$S(\nabla f(\beta),\lambda\alpha):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha)_{+},$
which is the soft thresholding operator.
@@ B.3 Theory
Let $\tilde{h}(\lambda):=([S(\nabla f(\beta),\lambda\alpha)]_{\mathcal{G},-0.5})_{\downarrow}$. Then taking $c=\tilde{h}(\lambda_{k+1})$ and $\phi=(1-\alpha)\lambda_{k+1}w$ as inputs for Algorithm A1 returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
@@
The proof is similar to that of Proposition 3.2. Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then,
@@
$\implies\text{cumsum}\Bigl{(}\bigl{(}[S(\nabla f(\beta),\lambda_{k+1}\alpha)]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}}\Bigr{)}$ $\prec,\mathbf{0}$
so that by the SGS subdifferential (Equation 7), all groups in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{g}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
@@
Suppose we have $\mathcal{B}\neq\emptyset$ after running the algorithm. Then, we have
$\text{cumsum}(\bar{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}\alpha v_{\mathcal{B}})\prec\mathbf{0}\implies\text{cumsum}\Bigl{(}\bigl{(}|\nabla f(\hat{\beta}(\lambda_{k+1}))|_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}\alpha v_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$ (15)
so that by the SGS subdifferential for non-zero groups (Equation 8), all variables in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{v}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{v}(\lambda_{k+1})$. ∎
@@
B.4 Path start derivation
@@
This is satisfied at the upper limit at $x=\tau\omega$, where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions, so that each variable within the same group is assigned the same value. Hence,
@@ Appendix C SLOPE Algorithm
Algorithm A1 Cumsum algorithm from [17]
@@
Appendix D Screening rule framework
@@ Appendix E Group-based OSCAR
This section presents an extension of the proposed screening rules for group-based OSCAR models. Firstly the OSCAR models are presented, and subsequently, the performance of the screening rules on synthetic data is presented.
E.1 Model definition
The Ordered Weighted $\ell_{1}$ (OWL) framework is defined as [35]
$\hat{\beta}=\operatorname*{arg\,min}_{\beta}\left\{\nabla f(\beta)+\lambda J_{\text{owl}}(\beta;v)\right\},$
where $J_{\text{owl}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, $|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$, and $v$ are non-negative non-increasing weights. SLOPE is a special case of OWL where the weights are taken to be the Benjamini-Hochberg critical values [4]. Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) [5] is a further special case of OWL (often referred to as OWL with linear decay) where for a variable $i$, the weights are taken to be $v_{i}=\sigma_{1}+\sigma_{2}(p-i)$, and $\sigma_{1},\sigma_{2}$ are to be set. In Bao et al. [3] they are set to $\sigma_{1}=d_{i}\|\mathbf{X}^{\top}y\|_{\infty},\sigma_{2}=\sigma_{1}/p$, where $d_{i}=i\times e^{-2}$.
Group OSCAR (gOSCAR) and Sparse-group OSCAR (SGO) are defined using the frameworks provided by gSLOPE [6] and SGS [14], respectively, but instead using the weights (for a variable $i$ and a group $g$)
@@ E.1 Model definition
with the weights visualised in Figure A3.
Figure A3: The SGO weights, $(v,w)$, shown for Figure 4 for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
@@ E.2 Results
Similar observations and conclusions made for the screening rules of gSLOPE and SGS are made for gOSCAR and SGO (Figures A4 - A8).
Figure A4 illustrates the effectiveness of the bi-level screening of SGO, similar to the effectiveness observed for SGS. Figures A5 - A6 showcase the efficiency of the screening rules on the proportion of the selected groups/variables. The screening rules are found to be effective across different data characteristics, with the running time of the models significantly decreasing (Figure A7). Similarly to SGS, KKT violations are more common compared to the ones observed for gOSCAR (Figure A8). This is due to the additional assumptions made at the additional screening layer of SGO and SGS. Similarly to Figure 6(a), the shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path.
Figure A4: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGO, shown for group and bi-level screening plotted as a function of the regularization path, applied to the synthetic data (Section 5.1). The data are generated under a linear model for $p=500,5000$. The results are averaged over 100 repetitions and 95% confidence intervals are shown (the SGO equivalent of Figure 1 (a)).
@@
Figure A7: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model. The results are averaged over $100$ repetitions, with standard errors shown (the OSCAR equivalent of Figure 4).
Figure A8: The proportion of KKT violations relative to the full input space, as a function of the regularization path. Group violations for gOSCAR and variable violations for SGS, fitted to linear models, averaged over all cases of $p$ and $\rho$ (the gOSCAR/SGO equivalent of Figure 6(a)).
@@ F.1 Computational information
The simulated experiments were executed on a high-performance computing cluster (x86-64 Linux GNU) and the real data analysis was conducted on a Apple Macbook Air (M1, 8GB).
F.2 Model Comparison
@@
Synthetic data.
For the linear model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-5}$ for gSLOPE and $10^{-8}$ for SGS (Table F.3). Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied). Of these model fits, there were no instances for gSLOPE where $\mathcal{E}$ was not a superset of $\mathcal{A}$. There was only one instance (out of the 98000) that this occurred for SGS, where $\mathcal{E}$ was missing a single variable contained in $\mathcal{A}$ (which had a non-screen fitted value of $\hat{\beta}=-0.004$).
For the logistic model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-8}$ for both gSLOPE and SGS (Table F.3.2). Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied). Of these model fits, there were no instances for gSLOPE or SGS where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
Real data.
In the real data analysis, the estimated coefficients with and without screening were very close to each other (with $\ell_{2}$ to be of order $10^{-8}$) for both SGS and gSLOPE (Table F.4). No instances occurred where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
@@ F.3 Additional results from the simulation study
Method Type Runtime (s)
Runtime (s) screen
@@
SGS Linear $735\pm 15$ $1830\pm 34$ $61\pm 1$ $84\pm 1$ $91\pm 1$ $21\pm 1$ $91\pm 3$ $708\pm 12$ $2\times 10^{-9}\pm 3\times 10^{-9}$
SGS Logistic $407\pm 2$ $859\pm 6$ $84\pm 1$ $107\pm 1$ $118\pm 1$ $22\pm 0.3$ $7\pm 0.2$ $51\pm 0.8$ $4\times 10^{-9}\pm 3\times 10^{-10}$
@@ F.3.1 Additional results for the linear model
Method $p$ Runtime (s)
Runtime (s) screen
@@
SGS $500$ $94\pm 1$ $133\pm 2$ $9\pm 0.2$ $11\pm 0.3$ $12\pm 0.3$ $3\pm 0.2$ $28\pm 1$ $74\pm 3$ $5\times 10^{-10}\pm 4\times 10^{-10}$
SGS $1625$ $416\pm 8$ $1129\pm 19$ $25\pm 1$ $37\pm 1$ $41\pm 1$ $12\pm 0.3$ $62\pm 7$ $511\pm 21$ $2\times 10^{-9}\pm 2\times 10^{-9}$
SGS $2750$ $639\pm 14$ $2137\pm 47$ $62\pm 2$ $82\pm 2$ $89\pm 3$ $20\pm 1$ $80\pm 6$ $791\pm 28$ $2\times 10^{-9}\pm 6\times 10^{-9}$
SGS $3875$ $939\pm 31$ $2862\pm 96$ $93\pm 3$ $124\pm 3$ $136\pm 4$ $30\pm 1$ $112\pm 8$ $1049\pm 34$ $5\times 10^{-9}\pm 1\times 10^{-8}$
SGS $5000$ $1586\pm 66$ $2891\pm 128$ $119\pm 4$ $164\pm 3$ $180\pm 4$ $39\pm 1$ $171\pm 11$ $1118\pm 37$ $2\times 10^{-9}\pm 4\times 10^{-9}$

### v2

@@ B.1 Penalty Weights
The penalty weights for SGS provide variable and group FDR-control simultaneously, under orthogonal designs (Feser and Evangelou, 2023). They are given by (where the indexing corresponds to the sorted variables/groups)
$v_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{1}{\alpha}F_{\mathcal{N}}^{-1}\left(1-\frac{q_{v}i}{2p}\right)-\frac{1}{3\alpha}(1-\alpha)a_{j}w_{j}\right\},\;i=1,\dots,p,$
$w_{i}^{\text{max}}=\max_{j=1,\dots,m}\left\{\frac{F_{\text{FN}}^{-1}(1-\frac{q_{g}i}{m})-\alpha\sum_{k\in\mathcal{G}_{j}}v_{k}}{(1-\alpha)p_{j}}\right\},\;i=1,\dots,m,$
where $F_{\chi_{p_{j}}}$ is the cumulative distribution function of a $\chi$ distribution with $p_{j}$ degrees of freedom, $F_{\mathcal{N}}$ is the cumulative distribution function of a folded Gaussian distribution, and $a_{j}$ is a quantity that requires estimation. The estimator $\hat{a}_{j}=\lfloor\alpha p_{j}\rfloor$ is proposed in Feser and Evangelou (2023). As with gSLOPE (Appendix A.1), a relaxtion is possible, giving the weights
@@
In the manuscript, as recommended by Feser and Evangelou (2023) under general settings, the SGS variable mean (Equation 11) and gSLOPE group mean (Equation 9) weights are used for all SGS numerical simulations.
@@ B.2 Derivation of Soft Thresholding Operator
    $\nabla_{i}f(\beta)\in[-\lambda\alpha v_{i},\lambda\alpha v_{i}]$: choose $y_{i}=\nabla_{i}f(\beta)/\lambda\alpha v_{i}$.
@@
$S(\nabla f(\beta),\lambda\alpha v):=\text{sign}(\nabla f(\beta))(|\nabla f(\beta)|-\lambda\alpha v)_{+},$
which is the soft thresholding operator. ∎
@@ B.3 Theory
Let $\tilde{h}(\lambda):=([S(\nabla f(\beta),\lambda\alpha v)]_{\mathcal{G},-0.5})_{\downarrow}$. Then taking $c=\tilde{h}(\lambda_{k+1})$ and $\phi=(1-\alpha)\lambda_{k+1}w$ as inputs for Algorithm A1 returns a superset $\mathcal{S}_{g}(\lambda_{k+1})$ of the active set $\mathcal{A}_{g}(\lambda_{k+1})$.
@@
The proof is similar to that of Proposition 3.2. Suppose we have $\mathcal{B}\neq\varnothing$ after running the algorithm. Then,
@@
$\implies\text{cumsum}\Bigl{(}\bigl{(}[S(\nabla f(\beta),\lambda_{k+1}\alpha v)]_{\mathcal{G},-0.5})_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}(1-\alpha)w_{\mathcal{B}}\Bigr{)}$ $\prec\mathbf{0},$
so that by the SGS subdifferential (Equation 7) all groups in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{g}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{g}(\lambda_{k+1})$. ∎
@@
Suppose we have $\mathcal{B}\neq\varnothing$ after running the algorithm. Then, we have
$\text{cumsum}(\bar{h}_{\mathcal{B}}(\lambda_{k+1})-\lambda_{k+1}\alpha v_{\mathcal{B}})\prec\mathbf{0}\implies\text{cumsum}\Bigl{(}\bigl{(}|\nabla f(\hat{\beta}(\lambda_{k+1}))|_{\downarrow}\bigr{)}_{\mathcal{B}}-\lambda_{k+1}\alpha v_{\mathcal{B}}\Bigr{)}\prec\mathbf{0},$
so that by the SGS subdifferential for non-zero groups (Equation 8) all variables in $\mathcal{B}$ are inactive. Hence, $\mathcal{S}_{v}(\lambda_{k+1})$ will contain the active set $\mathcal{A}_{v}(\lambda_{k+1})$. ∎
@@
B.4 KKT Checks
For SGS, the KKT conditions are first checked at the group-level for inactive groups (Appendix B.4.1). Further variable checks are performed for violating groups and variables in active groups (to check whether the variables should also be active) (Appendix B.4.2). The violating variables from these secondary variable checks are added back into $\mathcal{E}_{v}$.
B.4.1 Group Checks
A group violation occurs if the KKT conditions do not hold at the group-level (Equation 7). That is, a violation occurs if a group is discarded but
$\text{cumsum}\Bigl{(}\bigl{(}[\nabla f(\beta)+\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)]_{\mathcal{G}_{\mathcal{Z}},-0.5}\bigr{)}_{\downarrow}-\lambda(1-\alpha)w_{\mathcal{Z}}\Bigr{)}\succ\mathbf{0}.$
B.4.2 Variable Checks
For the set of variables in a violating group (from Appendix B.4.1), denoted $\mathcal{G}_{\mathcal{K}_{g}}$, a variable violation occurs if Equation 8 does not hold. That is, if
$\nabla_{\mathcal{G}_{\mathcal{K}_{g}}}f(\beta)\notin\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v_{\mathcal{G}_{\mathcal{K}_{g}}})\implies\text{cumsum}(|\nabla_{\mathcal{G}_{\mathcal{K}_{g}}}f(\beta)|-\lambda\alpha v_{\mathcal{G}_{\mathcal{K}_{g}}})\succ 0.$
B.4.3 Alternative KKT Checks
An alternative approach for SGS is to check the KKT conditions directly on the variables. The KKT conditions (Equation 6) can be rewritten as
$-\nabla f(\beta)-\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w)\in\lambda\alpha\partial J_{\text{slope}}(\beta;v).$
A KKT violation occurs the zero subdifferential conditions are not satisfied
$-\nabla f(\beta)-\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w)\notin\lambda\alpha\partial J_{\text{slope}}(\mathbf{0};v)$
$\implies$ $\text{cumsum}\left(\left|\nabla f(\beta)+\lambda(1-\alpha)\partial J_{\text{gslope}}(\beta;w)\right|_{\downarrow}-\lambda\alpha v\right)\succ\mathbf{0}.$
Now, the objective is to make the term inside the sorted absolute value operator as small as possible, given that the subdifferential term is unknown. To do this, a similar derivation as in Section B.2 can be used to determine that the term must be the soft thresholding operator, so that a violation occurs if
$\text{cumsum}\left(\left|S(\nabla f(\beta),\lambda(1-\alpha)\tau\omega)\right|_{\downarrow}-\lambda\alpha v\right)\succ\mathbf{0},$
where $\tau$ and $\omega$ are expanded vectors of the group sizes ($\sqrt{p_{g}})$ and penalty weights ($w_{g}$) to $p$ dimensions, so that each variable within the same group is assigned the same value. However, as we have had to approximate the unknown subdifferential term, this check is not exact. In practice, we found that this check was not stringent enough (due to the approximation), leading to violations being missed.
B.5 Path Start Proof
Proof of Proposition 4.4.
@@ B.5 Path Start Proof
This is satisfied at the upper limit at $x=\tau\omega$. Hence,
@@ Appendix C SLOPE ALGORITHM
Algorithm A1 SLOPE subdifferential algorithm from Larsson et al. (2020)
@@
Appendix D SCREEENING RULE FRAMEWORK
@@ Appendix E GROUP-BASED OSCAR
This section provides supplementary materials for extending the proposed screening rules to group-based OSCAR models.
E.1 Penalty Sequence
The gOSCAR and SGO weights are defined by (for a variable $i\in[p]$ and group $g\in[m]$) (Figure A3)
@@ E.1 Penalty Sequence
Figure A3: The SGO weights, $(v,w)$, for $p=500,m=100,q_{v}=0.05,q_{g}=0.05,\alpha=0.95$.
@@ E.2 Results
Observations and conclusions made for the screening rules of gSLOPE and SGS are also found to be true for gOSCAR and SGO (Figures A4 - A8).
Figure A4 illustrates the effectiveness of bi-selection of SGO, similar to the effectiveness observed for SGS. Figures A5 and A6 showcase the efficiency of the screening rules on the proportion of the selected groups/variables. The screening rules are found to be effective across different data characteristics, with the running time of the models significantly decreasing (Figure A7). KKT violations for SGO are more common compared to gOSCAR (Figure A8), due to the additional assumptions made at the second screening layer (as with SGS). Similar to Figure 7, the shape of the increasing number of KKT violations mirrors the log-linear shape of the regularization path.
Figure A4: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGO, shown for group and bi-level screening plotted as function of the regularization path, applied to the synthetic data (Section 6.1). The data are generated under a linear model for $p=500,5000$. The results are averaged over 100 repetitions and 95% confidence intervals are shown (the SGO equivalent of Figure 1).
@@
Figure A7: Runtime (in seconds) for fitting $50$ models along a path, shown for screening against no screening as a function of $p$, broken down into different correlation cases, for the linear model. The results are averaged over $100$ repetitions, with standard errors shown (the gOSCAR/SGO equivalent of Figure 4).
Figure A8: The proportion of KKT violations relative to the full input space, as a function of the regularization path. Group violations for gOSCAR and variable violations for SGS, under linear models, averaged over all cases of $p$ and $\rho$ (the gOSCAR/SGO equivalent of Figure 7).
@@ F.1 Computational Information
The simulated experiments were executed on a high-performance computing cluster (x86-64 Linux GNU) and the real data analysis was conducted on a Apple Macbook Air (M1, 8GB). Code for all simulations is available in the Supplementary Material. For all models, ATOS was used with the algorithmic parameters given in Table A1.
Table A1: Hyperparameters used for running ATOS in the synthetic and real data studies.
Parameter Synthetic Data Real Data
Max iterations 5000 10000
Backtracking 0.7 0.7
Max backtracking iterations 100 100
Convergence tolerance $10^{-5}$ $10^{-5}$
Convergence criteria $\|x-z\|_{2}$ $\|x-z\|_{2}$
Standardization $\ell_{2}$ $\ell_{2}$
Intercept Yes for linear Yes for linear
Warm starts Yes Yes
F.2 Solution Optimality
@@
Synthetic Data
For the linear model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-6}$ for gSLOPE and $10^{-9}$ for SGS (Table F.3). Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied). Of these model fits, there were no instances for gSLOPE where $\mathcal{E}$ was not a superset of $\mathcal{A}$. There was only one instance (out of the 98000) that this occurred for SGS, where $\mathcal{E}$ was missing a single variable contained in $\mathcal{A}$ (which had a non-screen fitted value of $\hat{\beta}=-0.004$).
For the logistic model, the maximum $\ell_{2}$ distances observed between the screened and non-screened fitted values were of order $10^{-8}$ for gSLOPE and $10^{-9}$ for SGS (Table F.3.2). Across the different cases, 98000 models were fit in total for each approach (excluding the models for $\lambda_{1}$, where no screening is applied). Of these model fits, there were no instances for gSLOPE or SGS where $\mathcal{E}$ was not a superset of $\mathcal{A}$.
Real Data
In the real data analysis, the estimated coefficients with and without screening were very close to each other for both SGS and gSLOPE (Table F.5). However, direct comparison is less meaningful here, as the models often failed to converge without screening, therefore not reaching the optimal solution.
@@ F.3 Additional Results from the Simulation Study
Method Type Runtime
Runtime screen (s)
@@
SGS Linear $735\pm 15$ $1830\pm 34$ $61\pm 1$ $84\pm 1$ $91\pm 1$ $91\pm 3$ $708\pm 12$ $2\times 10^{-9}\pm 3\times 10^{-9}$
SGS Logistic $407\pm 2$ $859\pm 6$ $84\pm 1$ $107\pm 1$ $118\pm 1$ $7\pm 0.2$ $51\pm 0.8$ $4\times 10^{-9}\pm 3\times 10^{-10}$
@@ F.3.1 Additional Results for the Linear Model
Figure A9: The proportion of groups/variables in $\mathcal{E},\mathcal{A}$, relative to the full input, for gSLOPE and SGS, as a function of the regularization path for the linear model with $p=2750,\rho=0.6,m=197$. The results are averaged over $100$ repetitions.
@@
Method $p$ Runtime
Runtime screen (s)
@@
SGS $500$ $94\pm 1$ $133\pm 2$ $9\pm 0.2$ $11\pm 0.3$ $12\pm 0.3$ $28\pm 1$ $74\pm 3$ $5\times 10^{-10}\pm 4\times 10^{-10}$
SGS $1625$ $416\pm 8$ $1129\pm 19$ $25\pm 1$ $37\pm 1$ $41\pm 1$ $62\pm 7$ $511\pm 21$ $2\times 10^{-9}\pm 2\times 10^{-9}$
SGS $2750$ $639\pm 14$ $2137\pm 47$ $62\pm 2$ $82\pm 2$ $89\pm 3$ $80\pm 6$ $791\pm 28$ $2\times 10^{-9}\pm 6\times 10^{-9}$
SGS $3875$ $939\pm 31$ $2862\pm 96$ $93\pm 3$ $124\pm 3$ $136\pm 4$ $112\pm 8$ $1049\pm 34$ $5\times 10^{-9}\pm 1\times 10^{-8}$
SGS $5000$ $1586\pm 66$ $2891\pm 128$ $119\pm 4$ $164\pm 3$ $180\pm 4$ $171\pm 11$ $1118\pm 37$ $2\times 10^{-9}\pm 4\times 10^{-9}$